OpenStudy (anonymous):

What is the volume of a cardboard box whose length, width, and height are 1/2 foot, 1/7 foot, and 1/3 foot, respectively? Would be amazing if you could help me!

OpenStudy (mathstudent55):

V = LWH \(V = \dfrac{1}{2} ft \cdot \dfrac{1}{7} ft \cdot \dfrac{1}{3} ft\) Multiply the 3 fractions together. The units are \(ft^3\). To multiply fractions together, multiply their numerators together, and multiply their denominators together.

OpenStudy (anonymous):

Holy moly thanks!

OpenStudy (mathstudent55):

You're welcome.